OpenStudy (anonymous):

rewrite using the distributive property and then simplify. 9(8+7)

OpenStudy (anonymous):

72 + 63 = 135

OpenStudy (anonymous):

Rewriting with the distributive property means distributing that nine to everything inside of the ( ) 9*8 + 9*7 72 + 63 135

OpenStudy (anonymous):

jimmyrep is that the correct anwser ?

OpenStudy (anonymous):

9(8+7)=9*8+9*7=72+63=135

OpenStudy (riley):

9(8) + 9(7) 72+ 63 =135 Just gotta break it up and solve. :)

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

thanks everyone